OpenStudy (anonymous):

What decimal is equivalent to seven and 3 over two hundred fifty? seven point one two seven point zero one two seven point three two five seven point three zero two five

OpenStudy (anonymous):

\[\LARGE 7\frac{3}{250} \quad \quad ?\]

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

whoa.. is that what you mean by "word discribing" ? lol

OpenStudy (anonymous):

hold on... I'll show you how to do it ;:)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\LARGE 7\frac{3}{250} =7+\frac{3}{250} =\frac{250\cdot 7}{250 }+\frac{3}{250}\] \[\LARGE \frac{1750+3}{250}= \frac{1753}{250}=7.012\] ... use calculator :)

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

hi msg me plz

OpenStudy (anonymous):

.. My pleasure. "hi msg me plz" ? O_O

OpenStudy (anonymous):

i ment hay

OpenStudy (anonymous):

is there anything you don't get in this question ? :)

OpenStudy (anonymous):

noe i get idd nw

OpenStudy (anonymous):

then I guess I'm leaving . ;) bye